# Instructions for *ACL Proceedings

### v1

@@ 6 Results
Table 3 shows average F1 scores across five runs (seeds 42, 0, 1, 2, 3) for all model configurations across our CS effectiveness dimensions. We compare BERT baselines, multi-task learning models, and dependency-based architectures (with CS embeddings). Dependency_matrix_6e consistently achieves the highest average F1 score (0.96), outperforming all the others. This performance gain is likely due to the use of a learnable dependency matrix that captures the interrelations between dimensions, allowing the model to exploit cross-dimensional dependencies, unlike models that treat the six dimensions independently. Statistical tests confirm the results differ significantly, especially between the best model, Dependency_m._6e, and all others (Appendix C). Figure 2 visualizes the dependency matrix learned by Dependency_m._6e. Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence. Emotional Appeal, for instance, is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84). These inter-dependencies help the model to improve its classification performance across dimensions. Both multi-task models, Multi-task_d. (0.94) and Multi-task_u. (0.92), outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided. Incorporating HS data in Bert_CS_HS (0.90) does not significantly improve performance over Bert_CS (0.91). Emotional Appeal remains the hardest dimension to classify, showing the most variation across models. In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_m._6e and Multi-task_d.), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues. Audience Adaptation scores above 0.99 in all configurations (Appendix E).
@@
Figure 2: Learned Dependency Matrix obtained from dep._m._6e on the combined dataset with seed 42.
@@
Cross-Domain Generalization. In the CONAN $\rightarrow$ Twitter setting, all models show a marked performance drop, with Dependency_m._6e performing best (0.66) alongside Dependency_m._3e and Multi-task_u. Emotional Appeal declines the most, suggesting a domain shift in how emotional content is conveyed. Clarity, Fairness, and Rebuttal scores are also low, reflecting the less structured nature of user-written CS. In the reverse setting, Twitter $\rightarrow$ CONAN, results are similar. Multi-task_d. achieves the highest avg. F1 score (0.66), followed closely by Dependency_m._6e (0.65). The remaining models obtain lower average F1 scores, indicating less robust generalization in this transfer setup. Models trained on expert-written CS seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55. In contrast, Audience Adaptation remains consistently high. In-Domain Evaluation. Training and testing on the same dataset (Appendix D) confirm that dependency-based models consistently perform best, further supporting the benefit of modeling interrelations between dimensions in classification.
7 Conclusion
As a first contribution of this paper, we release a novel linguistic resource that merges the CONAN and Twitter datasets for CS, manually annotated with six dimensions of CS effectiveness. As a second contribution, we present a framework for automatically classifying CS effectiveness across these dimensions. Results show that the dependency-based model outperforms BERT baselines with 0.96 F1 on both expert- and user-written CS. Multi-task models also surpass BERT with 0.94 F1, confirming that effectiveness dimensions are interdependent and benefit from explicit dependency modeling. Our architecture-agnostic framework improves BERT and generalizes to other transformers. A deep linguistic analysis highlights which dimensions need improvement for CS evaluation.
8 Limitation
Our framework for assessing CS effectiveness dimensions focuses on key aspects such as structural coherence, linguistic strategies, content strength, and rhetorical impact. Given the vast body of literature on counter-speech in social sciences, we focused on the most relevant studies, but acknowledge that further conceptualizations exist.
Another limitation concerns validating the effectiveness of CS with human users. Future work should assess whether CS aligned with our dimensions effectively influences discourse, fosters empathy toward targets, induces belief change in users, and reduces hateful content from the original authors.
Regarding the data, we acknowledge that our dataset is imbalanced, with a higher number of counter-speech examples generated by experts compared to those from online users, a common trait in counter-speech corpora due to the scarcity of high-quality, user-generated responses Susan et al. (2016); Mathew et al. (2019); Chung et al. (2019); Qian et al. (2019b). Additionally, some dimensions, particularly Audience Adaptation, exhibit class imbalance, which may affect model performance. To improve robustness, future studies should leverage larger and more balanced datasets.
Despite these limitations, our study introduces a new evaluation framework for counter-speech effectiveness and provides a model setup that achieves strong classification results. With this work, we aim to support and strengthen the community effort towards the refinement of counter-speech evaluation frameworks and the improvement of automatic approaches for counter-speech effectiveness assessment and generation.
9 Ethical Statement
We are aware that our study involves human exposure to hateful and potentially upsetting content. For this reason, we followed the mitigation strategies proposed by Vidgen et al. (2019) to protect annotators’ mental health. Participants gave their consent to voluntarily participate in the annotation, and were informed that they could withdraw from the study at any time. Moreover, each participant was allocated a small subset of the data (100 examples), and a generous time frame of about three weeks to mitigate prolonged exposure to potentially distressing language. We plan to share the data with the new annotations for research purposes only.
Appendix A Pairwise IAA Scores
Table 4 shows the results of the IAA between each pair of annotators on both CONAN and Twitter datasets, after the first round of annotations. Table 5 shows the pairwise IAA results after the second round. For binary metrics, we use Fleiss’s $\kappa$, while for multi-label metrics we use Krippendorff’s $\alpha$. For the audience adaptation dimension, we use Percent Agreement.
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.35 0.52 0.51 0.24 0.36 0.18
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.44 0.58 0.56 0.06 0.35 0.30
evidence 0.50 0.70 0.19 0.22 0.21 0.12
rebuttal 0.31 0.32 0.85 0.12 0.28 0.34
fairness 0.56 0.74 0.72 0.57 0.40 0.48
Table 4: First round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Dimension Ann1 & Ann2 Ann1 & Ann3 Ann2 & Ann3
Tw Co Tw Co Tw Co
emotional_appeal 0.88 0.60 0.42 0.85 0.55 0.48
audience_adaptation 1.00 1.00 0.99 1.00 0.99 1.00
clarity 0.88 0.76 0.86 0.79 0.73 0.71
evidence 0.78 0.78 0.85 0.77 0.64 0.62
rebuttal 1 0.72 0.48 0.90 0.48 0.72
fairness 0.92 0.90 0.95 0.78 0.94 0.70
Table 5: Second round IAA scores across dimensions for each annotator pair on Twitter (Tw) and CONAN (Co).
Appendix B Additional Experimental Details
In this section, we describe the model configurations used in our experiments. The goal of these experiments is to classify the effectiveness dimension scores for a given counter-speech response. More specifically, we use bert-base-uncased with 110M parameters. In all configurations, the input length is controlled by setting max_length to 128. The batch_size is 16, the random_state is set to 42, 0, 1, 2, and 3, and the learning_rate is 2e-5. All the inputs are tokenized by using BertTokenizer, and the process includes 3 training epochs (6 for two configurations). For performing the experiments, we concatenate both CONAN and Twitter datasets, and then we split it into train, validation, and test subsets, with a percentage of 70-10-20. We also perform cross-validation, by using a single dataset for training and validation, and the other for testing. The experiments were run on one A100 GPU.
We use the following loss functions:
  • Binary Cross-Entropy Loss (BCE): Used for binary classification tasks, it measures the difference between predicted probabilities and actual labels.
  • Cross-Entropy Loss (CE): Generalized for multi-class classification, CE extends BCE to multiple classes by computing the log loss over softmax-normalized probabilities.
  • Focal Loss: an extension of CE/BCE, focal loss addresses class imbalance by down-weighting easy samples and focusing on hard ones. It introduces a modulating factor to reduce the contribution of well-classified examples.
We also provide more details regarding our model contributions:
  • Multi-task_divided. We take the binary dimensions, and we train bert-base-uncased on them separately, then we combine the two losses (BCE and Focal Loss, for imbalanced datasets) to make the predictions. Then, we take the multi-label dimensions, we train them separately with Cross Entropy Loss, and, at the end, we combine all the losses, and we make the predictions on the test set. This configuration assumes that dimensions, for which the losses are added together, are somehow correlated.
  • Multi-task_united. Here, we take all the dimensions together, both binary and multi-label, and we train them with their respective loss functions (BCE, Focal Loss, and Cross-Entropy loss). Then we aggregate all the losses, and we predict the results for all of them. All predictions are done with the same Multi-task bert model. Again, the configuration assumes that dimensions are somehow correlated.
  • Dependency_learnable. All the dimensions are considered together, by training bert-base-uncased on them. Each dimension has its own loss function (BCE for emotional appeal, Focal Loss for audience adaptation, Cross Entropy Loss for all the multi-label dimensions), which are summed up at the end. In addition, there is a new parameter called dependency-matrix. Its rows and columns contain values between 0 and 1, that are weights accounting for the dependency between pairs of dimensions. Its weights are randomly initialized, and their best value is learned during the training phase. We run the model for 3 and 6 epochs.
We experiment with two types of input embeddings: (i) concatenated Hate Speech (HS) and Counter Speech (CS) embeddings to provide broader context and (ii) CS embeddings alone. The latter consistently yields slightly better performance. Consequently, we adopt CS-only embeddings for all configurations in the main part of the paper (Multi-task_d., Multi-task_u., Dependency_m._3e, and Dependency_m._6e). For completeness, Table 6 also reports the average results over five runs for models using concatenated CS+HS embeddings.
Model Emo. Aud. Clarity Evid. Rebut. Fair. AVG
Train: Combined Dataset
multitask_d. 0.90±0.02 0.99±0.00 0.92±0.01 0.90±0.01 0.91±0.03 0.91±0.03 0.92±0.01
multitask_u. 0.80±0.04 0.99±0.00 0.92±0.02 0.90±0.02 0.91±0.03 0.92±0.02 0.91±0.01
dep._m._3e 0.78±0.10 0.99±0.00 0.91±0.01 0.88±0.05 0.89±0.03 0.93±0.03 0.90±0.02
dep._m._6e 0.90±0.04 0.99±0.00 0.94±0.01 0.96±0.02 0.96±0.01 0.95±0.02 0.95*±0.01
Train: CONAN $\rightarrow$ Test: Twitter
multitask_d. 0.48±0.07 1.00±0.00 0.38±0.02 0.88±0.04 0.80±0.02 0.33±0.03 0.65±0.02
multitask_u. 0.51±0.06 1.00±0.00 0.40±0.03 0.88±0.04 0.79±0.02 0.35±0.05 0.65±0.01
dep._m._3e 0.56±0.02 1.00±0.00 0.39±0.02 0.90±0.02 0.79±0.02 0.34±0.05 0.66*±0.01
dep._m._6e 0.51±0.05 1.00±0.00 0.38±0.01 0.88±0.03 0.79±0.02 0.34±0.01 0.65±0.01
Train: Twitter $\rightarrow$ Test: CONAN
multitask_d. 0.40±0.17 0.99±0.00 0.49±0.24 0.43±0.01 0.60±0.05 0.86±0.13 0.63±0.04
multitask_u. 0.40±0.23 0.99±0.00 0.45±0.27 0.43±0.02 0.54±0.07 0.70±0.39 0.58±0.09
dep._m._3e 0.32±0.24 0.99±0.00 0.39±0.22 0.43±0.03 0.50±0.09 0.73±0.13 0.56±0.08
dep._m._6e 0.35±0.21 0.99±0.00 0.67±0.04 0.48±0.07 0.43±0.21 0.86±0.14 0.63*±0.07
Table 6: Effectiveness F1 scores (± std) for all models. Statistically significant results are marked with *.
@@ Appendix B Additional Experimental Details
Table 7: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS only embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Model Twitter Dataset CONAN Dataset
multi-task_d. 0.54 $\pm$ 0.19 1.00 $\pm$ 0.00 0.36 $\pm$ 0.24 0.97 $\pm$ 0.07 0.62 $\pm$ 0.25 0.53 $\pm$ 0.15 0.67 $\pm$ 0.05 0.94 $\pm$ 0.03 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.95 $\pm$ 0.02 0.94 $\pm$ 0.02 0.98 $\pm$ 0.03 0.95 $\pm$ 0.01
multi-task_u. 0.47 $\pm$ 0.19 1.00 $\pm$ 0.00 0.31 $\pm$ 0.12 0.97 $\pm$ 0.07 0.56 $\pm$ 0.19 0.36 $\pm$ 0.14 0.61 $\pm$ 0.07 0.89 $\pm$ 0.02 0.99 $\pm$ 0.00 0.91 $\pm$ 0.03 0.94 $\pm$ 0.02 0.96 $\pm$ 0.02 0.97 $\pm$ 0.01 0.92 $\pm$ 0.01
dep._m._3e 0.38 $\pm$ 0.14 1.00 $\pm$ 0.00 0.37 $\pm$ 0.07 0.97 $\pm$ 0.07 0.61 $\pm$ 0.19 0.48 $\pm$ 0.18 0.63 $\pm$ 0.06 0.84 $\pm$ 0.10 0.99 $\pm$ 0.00 0.96 $\pm$ 0.02 0.93 $\pm$ 0.03 0.94 $\pm$ 0.04 0.96 $\pm$ 0.04 0.94 $\pm$ 0.02
dep._m._6e 0.53 $\pm$ 0.16 1.00 $\pm$ 0.00 0.44 $\pm$ 0.20 0.97 $\pm$ 0.07 0.66 $\pm$ 0.21 0.56 $\pm$ 0.15 0.69 $\pm$ 0.03 0.95 $\pm$ 0.03 0.99 $\pm$ 0.00 0.98 $\pm$ 0.02 0.99 $\pm$ 0.01 0.98 $\pm$ 0.02 0.99 $\pm$ 0.02 0.98 $\pm$ 0.01
Table 8: F1 scores per effectiveness dimension on Twitter and CONAN datasets with CS+HS embeddings. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity of Arguments, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
Appendix C Statistical Significance Test
The dependency_matrix_6e model consistently achieves the highest average F1 scores across five runs, outperforming all other configurations. A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p < 0.001). To further assess whether dependency_m._6e is significantly better than the rest, we conduct pairwise t-tests between it and each of the other models. Results show that dependency_m._6e significantly outperforms bert_cs (p = 0.000114), bert_cs_hs (p = 0.000010), multi_task_d. (p = 0.003057), multi_task_u. (p = 0.000145), and dependency_m._3e (p = 0.000028) at the 1% significance level. These findings demonstrate that dependency_m._6e provides a statistically significant improvement over other baselines and variants.
Appendix D In-Data Results
Table 7 reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset, and for CONAN. The results are the average of the scores obtained from 5 runs with seeds of 42, 0, 1, 2, 3, with CS only embeddings, as the results in the main paper. Table 8 reports results with CS + HS embeddings. In both settings, we confirm the trend presented in the main paper, with the Dependency_matrix_6e achieving the best average F1 score. Furthermore, results show that models trained with only CS embeddings achieve slightly better scores than the ones trained on CS and HS embeddings. This can be due to the fact that adding more context, in this scenario, does not help the model to perform better, but it adds confusion. We also notice that training and testing on expert-written CS (CONAN) seems easier as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
Appendix E Error Analysis
In this section, we discuss some examples to better understand the models’ behaviors, in particular on the most difficult class (emotional appeal) and the one achieving almost perfect F1 score (audience adaptation).
@@ Appendix E Error Analysis
HS: "Islam is not a religion that professes peace and it does not make any good men. It condemns women and encourages delinquency. Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers."
CS: "Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men. For example, many women work as scientists, engineers, lawyers, etc." Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation. The CS provides counter-examples to the HS claims, using similarly complex and detailed arguments. This makes it easier for the classifier to recognize this dimension, as both speeches maintain a similar linguistic and argumentative structure. However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the high class imbalance. Further error analysis should investigate whether models are truly understanding adaptation strategies or simply leveraging surface-level cues.
@@
Twitter: "@user this isn’t funny at all. don’t make jokes of the situation when rene herself is affected by it. don’t get her dragged even more and delete this." This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns ("you").
CONAN: "I thought we’d stopped exporting our convicted criminals last century. Now you advocate exporting people without a trial or a conviction." This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument. The emotional appeal here is more subtle and indirect, contrasting with the explicit frustration in the Twitter example. A model trained solely on Twitter might miss these subtleties found in expert-written CS. To improve emotional appeal classification, domain adaptation or dataset augmentation could help models recognize diverse emotional expressions.

### v2

@@ VI Results
Table III shows average F1 scores across five runs for all model configurations across our counter-speech effectiveness dimensions. We compare BERT baselines, multi-task learning models, and dependency-based architectures (with CS embeddings). The dependency-based model Dependency_matrix_6e consistently achieves the highest F1 score (0.96), outperforming all the others. This performance gain is likely due to the use of a learnable dependency matrix that captures the interrelations between dimensions, allowing the model to exploit cross-dimensional dependencies, unlike models that treat the six dimensions independently. Statistical tests confirm that the results are significantly different. A one-way ANOVA test, conducted on the average F1 scores across models, confirms a statistically significant difference in performance (F = 12.92, p << 0.001), especially between the best model, Dependency_matrix_6e, and all others (pairwise t-tests between it and each of the other models show statistically significant differences at the 1% significance level).
Figure 2 shows the dependency matrix learned by Dependency_matrix_6e. Rows represent influenced tasks, columns represent influencing tasks, with higher values indicating stronger influence (diagonal values are zero as dimensions do not influence themselves). We can see that Emotional Appeal is influenced by Fairness (0.63) and Rebuttal (0.65), while Audience Adaptation is shaped by Clarity (0.98), Rebuttal (0.96), and Fairness (0.84). Clarity is strongly influenced by Rebuttal (0.95) and Emotional Appeal (0.83); Evidence is influenced by Audience Adaptation (0.98); Fairness is influenced by Emotional Appeal (0.80) and Audience Adaptation (0.95), and Rebuttal by Emotional Appeal (0.7). These inter-dependencies help the model to improve its classification performance across dimensions. Both multi-task models, Multi-task_divided (0.94) and Multi-task_united (0.92), also outperform the BERT baselines, suggesting that the multi-task framework is more effective, with only a slight performance difference between united and divided. Fine-tuned BERT models remain competitive: Bert_CS (0.91) slightly outperforms Bert_CS_HS (0.90). This suggests that incorporating HS data does not significantly improve classification performance.
Emotional Appeal remains the hardest dimension to classify, showing the most variation across models, ranging from 0.67 to 0.92 F1 for the combined dataset. In this dimension, dependency-based and multi-task models perform best (up to 0.92 for Dependency_matrix_6e and Multi-task_divided), while BERT baselines lag behind (0.67 and 0.62), suggesting that richer architectures better capture emotional cues. Audience Adaptation scores above 0.99 in all configurations. Although this can be due to class imbalance, we address this concern using focal loss during training, which focuses the training on hard, misclassified examples and reduces the impact of easy, well-classified examples. By contrast, Clarity and Evidence show greater variation, pointing to their higher classification difficulty.
@@
Figure 2: Learned Dependency Matrix obtained from Dependency_matrix_6e on the combined dataset with seed 42.
@@
Cross-Domain Generalization. To assess model generalization, we perform cross-domain evaluation. In the first setting, models are trained on CONAN (expert-written CS) and tested on Twitter Dataset (user-generated CS). All models show a marked performance drop, with Dependency_matrix_6e performing best (0.66) alongside Dependency_matrix_3e and Multi-task_united. Emotional Appeal declines the most, suggesting a domain shift in how emotional content is conveyed. Clarity, Fairness, and Rebuttal scores are also low, reflecting the less structured nature of user-written CS. In the reverse setting, where models are trained on Twitter Dataset (user-written CS) and tested on CONAN (expert-generated CS), results are similar. Multi-task_divided achieves the highest F1 score (0.66), followed closely by Dependency_matrix_6e (0.65), which remains highly competitive. The remaining models obtain lower average F1 scores, indicating less robust generalization in this transfer setup. Models trained on expert-written CS (Twitter $\rightarrow$ CONAN) seem more adaptable, but Emotional Appeal remains challenging across domains, ranging from 0.14 to 0.55. In contrast, Audience Adaptation remains consistently high, showing robustness across domains, while Fairness remains strong only on expert-written data. Overall, our models generalize reasonably well, achieving a good F1 score in these cross-dataset tests. This suggests that models learn meaningful, transferable patterns beyond dataset-specific characteristics. Given the variability in emotional appeal, domain adaptation strategies or dataset augmentation could help models better capture diverse emotional expressions across different counter-speech sources.
TABLE IV: F1 scores per effectiveness dimension on Twitter and CONAN datasets. AVG: macro-average. Emo.: Emotional Appeal, Aud.: Audience Adaptation, Clarity: Clarity, Evid.: Evidence, Rebut.: Rebuttal, Fair.: Fairness.
@@
In-Domain Evaluation. Table IV reports F1 classification results of model configurations, with training and testing on the same dataset for the Twitter Dataset and for CONAN. Training and testing on the same dataset confirm that dependency-based models consistently perform best across effectiveness dimensions, with Dependency_matrix_6e again achieving the best average F1 score, further supporting the benefit of modeling interrelations between dimensions in classification tasks. We also notice that training and testing on expert-written CS (CONAN) seem easier, as it achieves higher average F1 scores, compared to training and testing on the Twitter dataset.
The results across all three evaluation settings highlight key insights into the classification of counter-speech effectiveness dimensions. While emotional appeal presents unique challenges due to its subjectivity, the other dimensions (audience adaptation, clarity, evidence, rebuttal, and fairness) also exhibit distinct patterns that provide valuable directions for model improvement.
VII Error Analysis and Discussion
In this section, we discuss some examples to better understand the models’ behaviors on two dimensions, Emotional Appeal (most difficult to classify) and Audience Adaptation (highest F1 scores), and we address some limitations.
@@ VII Error Analysis and Discussion
HS: ”Islam is not a religion that professes peace and it does not make any good men. It condemns women and encourages delinquency. Lawbreakers are not all of a certain race or religion, but only one race and religion make lawbreakers.”
CS: ”Today, in many Muslim-majority communities, women receive equal, even higher in some places, treatment as men. For example, many women work as scientists, engineers, lawyers, etc.” Both speeches aim to persuade a specific group by matching linguistic complexity, thus ensuring audience adaptation. The CS provides counter-examples to the HS claims, using similarly complex and detailed arguments. This makes it easier for the classifier to recognize this dimension, as both speeches maintain a similar linguistic and argumentative structure. However, this near-perfect performance could reflect data bias, suggesting that audience adaptation might be too easily detected, due to the class imbalance. Further error analysis should investigate whether models are truly understanding adaptation strategies or simply leveraging surface-level cues.
@@
Twitter: ”@user this isn’t funny at all. don’t make jokes of the situation when rene herself is affected by it. don’t get her dragged even more and delete this.” This CS is direct and uses imperatives to express frustration and urgency, relying on internet culture and social cues like tagging someone and using second-person pronouns (”you”).
CONAN: ”I thought we’d stopped exporting our convicted criminals last century. Now you advocate exporting people without a trial or a conviction.” This example uses sarcasm, rhetorical questioning, and historical references to evoke moral outrage and challenge the opposing argument. The emotional appeal here is more subtle and indirect, contrasting with the explicit frustration in the Twitter example. A model trained solely on Twitter might miss these subtleties found in expert-written CS. To improve emotional appeal classification, domain adaptation or dataset augmentation could help models recognize diverse emotional expressions.
Model Limitations. We use BERT-based classifiers throughout to ensure that improvements stem from our multi-task and dependency-based framework rather than the underlying model. Our goal was not to benchmark architectures but to evaluate the framework itself. For example, while Large Language Models (LLMs) are promising, and can achieve good results, they are optimized for generation rather than classification tasks. Evaluating LLMs would also require a separate experimental design, which is beyond the scope of this paper, and we leave these comparisons as future work.
VIII Conclusion
As a first contribution of this paper, we release a novel linguistic resource to the community, enriching the CONAN and Twitter dataset with manual annotations of six fine-grained dimensions of counter-speech effectiveness: Emotional Appeal, Audience Adaptation, Clarity, Evidence, Rebuttal, and Fairness. This new annotation layer provides a richer, more diverse foundation for research on CS evaluation, considering content from both expert- and user-generated sources.
As a second contribution, we introduce a framework for automatically classifying counter-speech effectiveness across these dimensions. Our approach is architecture-agnostic and compatible with a range of transformer backbones. Experimental results demonstrate that a dependency-based model, which explicitly models inter-dependencies among the dimensions, significantly outperforms standard BERT baselines, achieving 0.96 F1 score on the combined dataset. This includes strong performance on both expert- (CONAN) and user-written (Twitter) CS. In addition, our multi-task learning models also surpass BERT with 0.94 F1 score, reinforcing the hypothesis that the six dimensions are interrelated and can be jointly learned more effectively. Our cross-domain evaluation reveals substantial domain shifts between CONAN and Twitter, especially in Emotional Appeal, highlighting the challenges of generalization and the need for domain-adaptive modeling techniques. Despite this, our framework demonstrates robust improvements over BERT baselines across domains, suggesting its effectiveness can generalize beyond specific model architectures.
Finally, we conduct a detailed linguistic analysis of model performance across dimensions, revealing which aspects of CS effectiveness remain the most difficult to capture. This analysis can inform the design of future CS evaluation systems and guide annotation efforts where additional human oversight is most valuable.
Acknowledgment
This work has been partially supported by the French government, through the 3IA Côte d’Azur investments in the project managed by the National Research Agency (ANR) with the reference number ANR-23-IACL-0001.